Vector Ā has magnitude 6 units and is in the direction of the + x axis. Vector Ā has magnitude 4 units and lies in the xy-plane, making an angle of 30 degree with the + x-axis. Find the vector product C= Ā × B vector
1
Expert's answer
2019-06-26T08:59:26-0400
"\bold{C}=\bold{A}×\bold{B}=\begin{vmatrix}
 \bold{i} & \bold{j} & \bold{k} \\
6 & 0 & 0 \\
 4\cos{30} & 4\sin{30} & 0 
\end{vmatrix}"

"\bold{C}=(6(4)\sin{30})\bold{k}=12\bold{k}"